OpenStudy (blurryface):

Find the perimeter of parallelogram ABCD with vertices A(–5, 6), B(2, 6), C(1, –2), and D(8, –2). I got the answer 28 is this correct?

OpenStudy (danjs):

what are the lengths of the sides?

OpenStudy (blurryface):

unknown

OpenStudy (danjs):

right, you calculate those..

OpenStudy (blurryface):

I did but I'm not sure if I did it correctly....

OpenStudy (eliesaab):

AB=?

OpenStudy (blurryface):

9

OpenStudy (eliesaab):

AB=7

OpenStudy (eliesaab):

CD=7

OpenStudy (eliesaab):

AC=?

OpenStudy (blurryface):

5

OpenStudy (eliesaab):

AC=10

OpenStudy (eliesaab):

10+10 +7+7=34

OpenStudy (blurryface):

ugh I did it wrong could you explain how you got AB and CD

OpenStudy (eliesaab):

A and B have the same y-coordinates AB=5+2

OpenStudy (eliesaab):

Take the point M=(1,6). The triangle AMC is right triangle AM=6, MN=8 AC^2= 36+64=100 AC=10

OpenStudy (eliesaab):

You can also use the distance formula between two point (a,b) and (c,d) \[ d=\sqrt{(a-c)^2 +(b-d)^2}\]

OpenStudy (eliesaab):

This will allow you to compute all the sides without using my shortcuts

OpenStudy (eliesaab):

Good night. I have to go to sleep

OpenStudy (danjs):



OpenStudy (danjs):

The horizontal lengths are easy to see the lengths, from x=-5 to x=2, 7 units same for the bottom piece, x=8 to x=1 , 7 units The slanted parts you can make right triangles and use the pythagorean theorem...

OpenStudy (danjs):

Sides 8 and 6